OpenStudy (shes.radical):

I already got the answer but can someone please check it? The following table shows the length and width of two rectangles: Length Width Rectangle A 2x + 4 x − 1 Rectangle B 3x + 1 2x − 3 Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? (a) 4x-10 <-- My Answer (b) 4x-14 (c) 16x+18 (d) 16x+14

OpenStudy (anonymous):

Perimeter of B: \[2(3x+1)+2(2x-3) = 10x - 4\] Perimeter of A: \[2(2x+4) + 2(x-1) = 6x + 6\] \[(10x - 4) - (6x + 6) = 4x - 10\] You are correct.

OpenStudy (shes.radical):

thanks!